Fermi National Laboratory
 

Accelerator Update for October 18, 2007 - October 19, 2007

 

Thursday October 18

The day shift began with the SciBooNE / MiniBooNE experiments taking beam when available, with the Neutron Therapy Facility (NTF) treating patients, with the Antiproton Source (Pbar) in access, and with the Main Injector (MI) and Tevatron (TeV) conducting studies.

Crew Chief Summary: During the Pbar access, technicians found two quadrupole magnets with LCW leaks (A1Q11 and D1Q13). Both magnets were repaired by 2:51 PM. Linac experts accessed their tunnel and worked on a kicker (KPS4A).

During the evening shift Pbar conducted studies. Work on PAK3 held off TeV studies. TeV suffered a few aborts due to false anode supply fault indications; operators replaced a control module at TeV sector F0 to solve the problem. A Booster quadrupole power supply caused some down time.

Friday October 19

The midnight shift began with Pbar stacking, with experts tuning the TeV, and with Sci/MiniBooNE taking beam.

Crew Chief Summary: A Recycler expert completed his studies at 4:18 AM. At 5:06 AM the Sci/MiniBooNE auto tune program began to wander. Operators manually corrected the problem several times and determined that it wasn’t an autotune problem; the beam seemed to be moving at VP869, but might have just been a readback problem. TeV experts accelerated protons to 980 GeV.

The Plans for Friday

The plans for today are to transfer antiprotons from Pbar to the Recycler. Accelerator commissioning continues.

Machine Summaries

Linac: Current status: output is ~ 35 mA. Weekly Statistics: 94% availability with 9.53 total hours of down time. NTF Treatment Schedule: calibration and treatment is on Monday, Thursday, and Fridays from 7 AM to 12 PM. Major Downtime: a new LRF2 6544 tube failed with not many hours of use (this is a very rare occurrence.) Work In Progress or Completed: switched the low energy Linac to individual station RF triggers, did a tune up of LRF5 (tuned on the 9” trombone, retuned cavity coupling loop to pre-shutdown position, achieved optimal beam loading), and continue to tune to reduce energy variations.

Booster: Experts have worked on extraction timing and orbits. They have worked on skew quadrupoles and ramps. They have tuned on vertical orbits. The horizontal orbits need more work, but they must wait for the Linac energy to steady. They have done RF measurements and tune-ups. Their intensity is at 6-7 turns and they have an efficiency of ~ 84%.

Main Injector: They have established an 8 GeV beam. They have completed MI-8 line lattice measurements and matching; they need to analyze the results of this before measuring again. They have established a 120 GeV beam and smoothed orbits, but this still needs a little more work. They have completed aperture scans at MI-10, MI-20, MI-40, and MI-52. They have decided not to do quadrupole moves. They are 50% done with a coalescing tune up and a slip stacking tune up. They have tuned up the extraction positions and the abort line, but still need to work on NuMI. Parasitic Measurements: commission new damper, collimator commissioning (MI-8 and MI-30), ring aperture scans with new software program, general tune up. Current Problems: They have an LCW leak on the MI-303 collimator (this will take 12 hours to fix), the MI-605 BPM is broken and might require an access to fix, there are leaks in the compressed air lines in the tunnel.

Antiproton Source: The change they made to their Pbar Debuncher momentum cooling system didn’t work. They reverted back to the way it was. They phased various systems. Stacking went very well. They are investigating the drop in the stacking rate that occurred on the midnight shift.

Recycler: They are ready for antiprotons. They still want protons for studies. The Electron Cooling experts are worried about the energy match.

Tevatron: Here is a list of what the experts have accomplished: established circulating beam, set orbit tunes at 150 GeV, reverse injection tune up, and ramped to flattop and smoothed the beam. Their present status: they have good orbits, but they have to make sure they have made the correct compromises between corrector strength and positions at IPs. The reverse injection works on the new helix, but they need to verify that the new helix at injection is exactly what they want.

FESS: They completed the wire splice for the 95 Pbar LCW pump #8 and would like to bring it on line. They will then need to check pumps 6 and 7. The Wilson Hall fire pump test will occur in about two weeks. They will be checking the SwitchYard strainers today.

Switchyard 120: Shutdown continues

NuMI: Shutdown continues.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/19/2007    email Fermilab